OpenStudy (anonymous):

An audio cable is 14.9 dm long. How long is the cable in meters?

OpenStudy (anonymous):

1dm=10cm 10dm= 1m can you do from now?

OpenStudy (anonymous):

(1m/m)=(10dm/14.9dm) time and flip =14.9/10 =1.49meters

OpenStudy (anonymous):

thanks again=)